OpenStudy (anonymous):

Admission to the movies costs $8 for adults and $4 for kids. The total bill for your entire extended family was $72. This scenario can be modeled by the equation 8a + 4k = 72. Which is not a possible combination of adults and kids? 3 adults and 12 kids 5 adults and 6 kids 7 adults and 4 kids 8 adults and 2 kids

OpenStudy (anonymous):

just substitute the posibilities in the equation and check

OpenStudy (anonymous):

for example the 1º posibility: a=3, k=12 8*3+4*12=24+48=72 OK do the same with the rest

OpenStudy (anonymous):

thanks